Please help me with this problem:
(1/4)(8m-3) = (4/7)(3m-4)+9
Solving for m, of course.
I know I have to multiply by the common denominator, but I forgot all of which I multiply.
Your help would be greatly appreciated!

Answer by CharlesG2(834)

" Please help me with this problem:
(1/4)(8m-3) = (4/7)(3m-4)+9
Solving for m, of course.
I know I have to multiply by the common denominator, but I forgot all of which I multiply.
Your help would be greatly appreciated! "
(1/4)(8m-3) = (4/7)(3m-4)+9
2m - 3/4 = (4/7)(3m) -4(4/7) +9
2m - 3/4 = (12m)/7 -16/7 +9
14m - 21/4 = 12m - 16 + 63 (multiplied all terms by 7)
2m = 21/4 - 16 + 63
2m = 21/4 + 47
m = 21/8 + 47/2
m = 21/8 + (47*4)/(2*4)
m = 21/8 + 188/8
m = 209/8
m = 26 1/8